The Hamiltonian of a system is given in the form of the equation H = 1/2 (1/q^2 + p^2 q^4)

Write down the equation of motion for q !


1
Expert's answer
2021-07-05T16:13:45-0400

We know that

"-\dot{p}=\frac{dH}{dq},\dot{q}=\frac{dH}{dp}"

"-\dot{p}=\frac{d}{dq}(\frac{1}{2q^2}+\frac{p^2q^4}{2})"

"\dot{p}=\frac{1}{q^3}-p^2q^3"

"\dot{q}=\frac{dH}{dp}"

"\dot{q}=\frac{d}{dp}(\frac{1}{2q^2}+\frac{p^2q^4}{2})"

"\dot{q}=pq^4\rightarrow{1}"

Equation (1) is called eqution of motion q